2|x-3|=10 what is x?
can you explain this question so i can understand it?

Answer by ankor@dixie-net.com(22740)

2|x-3| = 10
divide both sides by 2
|x - 3| = 5
Since it is in an absolute, x can be positive and x can be negative
Two solutions
x - 3 = 5
x = 5 + 3
x = 8
and
x - 3 = -5
x = -5 + 3
x = -2
;
;
Check both solutions in original problem
x = 8
2|8-3|= 10
2(5) = 10
and
x = -2
2|-2-3| = 10
2|-5| = 10
Any value coming out of the absolute is positive, therefore
2(5) = 10